A motorcycle slows down on a sharp horizontal turn, going from 30.0 m/s to 10.0 m/s in 15.0 s . The radius of the curve is 166 m. Assuming the motorcycle slowing down during this time at the constant rate, find the magnitude of the total acceleration in m/s2 at the moment the motorcycle speed reaches 10.0 m/s. 


1
Expert's answer
2021-02-28T07:39:28-0500

"u=30m\/s"

"v=10m\/s"

"t=15 s"

"r=166m"

"\Rightarrow v=u+a_tt"

"\Rightarrow a_t=-\dfrac{4}{3}m/s^2"


Centripetal acceleration, "a_c=\dfrac{v^2}{r}"

"\Rightarrow a_c=\dfrac{(10)^2}{166}"

"\Rightarrow a_c=0.60m/s^2"


Total Accceleration, "a_{total}=\sqrt{a_c^2+a_t^2}"

"\Rightarrow a_{total}=\sqrt{(0.60)^2+(1.33)^2}"

"\Rightarrow a_{total}=1.45m/s^2"